ensen Boat Works is an all equity firm that has 340,000 shares of stock outstanding. The company is in the process of borrowing $4 million at 8% interest to repurchase 80,000 shares of the outstanding stock. What is the value of this firm if you ignore taxes?
Answer #1

Answer:

We need to find share price of the company.

The company is in the process of borrowing $4 million at 8% interest to repurchase 80,000 shares of the outstanding stock.

Hence:

Market price of share = 4000000 / 80000 = $50

It is an all equity company.

Hence:

Value of this firm = 340000 * 50 = $17,000,000

Value of this firm = $17,000,000 = $17 million